OpenStudy (anonymous):

Find the length of the hypotenuse of a right triangle whose legs are 3 and sqrt(2) a) 2 sqrt(2) b) sqrt(11) c) sqrt(13)

jimthompson5910 (jim_thompson5910):

a^2 + b^2 = c^2 (3)^2 + (sqrt(2))^2 = c^2 9 + 2 = c^2 .. .. .. c = ??

OpenStudy (anonymous):

is it B

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

you're welcome